OpenStudy (loser66):

Please check, Find annihilator of \[F(x)= e^{-3x}(2sinx +7cosx)\]

OpenStudy (loser66):

since sinx, and cosx are linearly independent, I break them out and got the annihilator for each of them is A(D) = (D^2+6D +10) so, the final answer is \(A_total (D) = (D^2 +6D +10)^2\) but not sure about ^2 Please help

OpenStudy (helder_edwin):

what subject is this for?

OpenStudy (loser66):

differential equation

OpenStudy (loser66):

@jdoe0001

OpenStudy (loser66):

no one help me???? why??

OpenStudy (loser66):

@hartnn @phi @terenzreignz

OpenStudy (loser66):

@hba @austinL

terenzreignz (terenzreignz):

I don't even know what a bloody annihilator is... googling... brb :>

OpenStudy (austinl):

^ I have the same issue.... what the deuce is an annihilator?

OpenStudy (loser66):

do you want to see my book? it's differential equation, I sssswwweaar

terenzreignz (terenzreignz):

Have I ever been known to doubt you L66? :> ...of course I don't :D

OpenStudy (loser66):

@austinL do you know linear operator in dff?

OpenStudy (austinl):

It sounds vaguely familiar....

OpenStudy (loser66):

ok, let me ask my prof, then. thanks anyway.

OpenStudy (austinl):

I will ask my professor in just a moment... I will see what he has to say.

terenzreignz (terenzreignz):

wait, I think I may have gotten something...

OpenStudy (loser66):

this what my book says

terenzreignz (terenzreignz):

Okay, I did get the same answer as you did... pretty much, and I share your suspicion that maybe the ^2 isn't necessary... so you know what, let's just differentiate the blasted thing XD

OpenStudy (loser66):

@terenzreignz you mean no need to take square?

terenzreignz (terenzreignz):

Let's find out. After all, both terms have the same annihilator, so why bother doing it twice? But let's try to find out the hard way, see if it works...\[\Large (D^2+6D +10)\left[e^{-3x}(2\sin x + 7 \cos x)\right]\]

OpenStudy (loser66):

because sinx, and cos x are linearly independent

terenzreignz (terenzreignz):

Okay, let's do the first term... \[\Large (D^2+6D+10)2e^{-3x}\sin(x) \]

OpenStudy (loser66):

so, we try both way, OK? you do just one, I do ^2 , or else?

terenzreignz (terenzreignz):

Let's do just 1, and it it doesn't work, then it's probably ^2. Don't kill yourself, LOL

OpenStudy (loser66):

oh no!! for sure, friend, you see, just at the first line, you pick 2 as coefficient, I can conclude that 2 \(\neq\)7 of cos.

OpenStudy (loser66):

therefore, they cannot be the same. hehehe... is my logic reasonable?

terenzreignz (terenzreignz):

Those 2 and 7 are irrelevant, they are simply constants multiplied to the function and will of course have no effect on the annihilator (since if the function is annihilated [to zero] that constant multiplied to it can do nothing)

OpenStudy (loser66):

how can it be irrelevant?? because at the end up , you must let them =to calculate A, B the coefficient of cos and sin from partial solution.right?

terenzreignz (terenzreignz):

Just hang on :>\[\Large (D^2+6D+10)2e^{-3x}\sin(x)\] amounts to \[-4e^{-3x}(3\cos(x) - 4\sin(x) )+12e^{-3x}(\cos(x) - 3\sin(x))+20e^{-3x}\sin(x) \] which is zero...which makes sense, because the differential operator was the annihilator of \(\large 2e^{-3x}\sin x\) right?

OpenStudy (loser66):

yes

terenzreignz (terenzreignz):

Now, two things, first, if I'm your friend, acknowledge that I have a name, and it's TJ ^_^ Secondly, let me demonstrate to you the irrelevance of constants. Suppose we have a function f(x) and it has an annihilator a(D) You following? :>

OpenStudy (loser66):

I follow both (your 2 things) hehehe..TJ

terenzreignz (terenzreignz):

Well, anyway, it only follows that \[\large a(D)f(x)=0\] But what if f(x) is multiplied by some constant k? Would it matter? \[\Large a(D) kf(x) = ka(D)f(x) =k(0) = 0\]The constant factor won't matter. Got it? :>

OpenStudy (loser66):

just in the case you need to find out the annihilator of f (x). When putting it into the whole problem to find \(y_c~~and~~y_p\) it is not irrelevant, am I right ?TJ

terenzreignz (terenzreignz):

Get to that later. For now, we know that \[\Large (D^2+6D+10)2e^{-3x}\sin(x)=0\] But what about \[\Large (D^2+6D+10)7e^{-3x}\cos(x)\qquad \color{red}?\]

OpenStudy (loser66):

=0 , too

OpenStudy (loser66):

ah, we have \(A_1(D)*A_2(D) \) is annihilator of \(F_1(x) +F_2(x)\) right ? TJ

terenzreignz (terenzreignz):

Well yes, it should. Now, that can only mean that \[\Large (D^2+6D+10)[2e^{-3x}\sin(x)+7e^{-3x}\cos(x)]=0\]\[\Large =\Large (D^2+6D+10)e^{-3x}(2\sin(x)++7\cos(x))=0\]

terenzreignz (terenzreignz):

Hold on a sec, let's have another theoretical approach... what if we have \(f_1\) and \(f_2\) as functions of x with *the same* annihilator a(D) ? Then... \[\Large a(D)[f_1(x)+f_2(x)]= a(D)f_1(x)+ a(D) f_2(x)\]\[\Large = 0+0\]since a(D) was an annihilator to both of them, individually. \[\Large = 0\]and therefore, a(D) is the annihilator to their sum as well.

terenzreignz (terenzreignz):

Such is the case with \(\large 2e^{-3x}\sin(x)\) and \(\large 7e^{-3x}\cos(x)\) They do, in fact, have the same annihilator.

terenzreignz (terenzreignz):

Everything cool? :>

OpenStudy (loser66):

I still have another question, can I ask you here? the same topic

terenzreignz (terenzreignz):

But this one question is closed, then? Not bad for a kid that didn't know what annihilators were just half an hour ago, right? ;) That said, that was LUCK. So I can't guarantee your second question, but hit me up anyway :P

OpenStudy (loser66):

ok, I will post in a new post. Thanks a lot TJ

terenzreignz (terenzreignz):

You do that, we'll have a look see :>